are 3x-6y=12 and y=x+4/2 parallel?
Answer by mananth(16946)

3x-6y=12
convert equation to slope intercept form
6y=3x-12
/6
y=(1/2)x-2
m= slope = 1/2

and y=(x+4)/2
multiply by 2
2y= x+4
/2
y=(1/2)x+2
m= slope = 1/2
the slopes equal
so the lines parallel